A proton is moving at 425 m/s. (a) How much work must be done on it to stop it?(answer in joules) (A proton has a mass of 1.67×10−27 kg.) (b) Assume the net braking force acting on it has magnitude 8.01×10−16 N and is directed opposite to its initial velocity. Over what distance must the force be applied?(in meters) Watch your negative signs in this problem.
1
Expert's answer
2011-03-09T06:52:28-0500
(a) While moving with a speed of 425 m/s proton has the following kinetic energy:
K = mv2/2 = 1.67×10−27* 4252/2 = 1.51×10−22 J
Required work is equal to this energy.

(b) S = A/F = 1.51×10−22 / 8.01×10−16 N = 0.189×10−6 m.